OpenStudy (anonymous):

What are the vertex and the axis of symmetry of the equation? y = –2x2 + 8x – 18 vertex: ( 2, –10) axis of symmetry: x = 2 vertex: (2, –10) axis of symmetry: x = –10 vertex: (–2, –10) axis of symmetry: x = –2 vertex: (–2, 10) axis of symmetry: y = –2

OpenStudy (helder_edwin):

u have to complete the square: \[ \large y=-2x^2+8x-18 \] \[ \large y+18\color{red}{-8}=-2(x^2-4x\color{red}{+4}) \] \[ \large y+10=-2(x-2)^2 \]

OpenStudy (anonymous):

So , its the last one , rite ?

OpenStudy (anonymous):

& thank you for showing me how to do it (: